# Hyperbolic nonlinear Schrödinger equations on $\mathbb{R}\times\mathbb{T}$

### v2

@@ 1. Introduction
$\begin{cases}i\partial_{t}u+\Box u=\pm|u|^{2k}u,\\ u(t,\mathbf{x})|_{t=0}=u_{0}(\mathbf{x}),\end{cases}\quad(t,\mathbf{x})\in\mathbb{R}\times\mathcal{M},$ (1.1)
@@
$\|e^{it\partial_{x}\partial_{y}}f\|_{L^{q}_{t}L^{p}_{x,y}(\mathbb{R}\times\mathbb{R}^{2})}\lesssim\|f\|_{L^{2}(\mathbb{R}^{2})}$ (1.3)
@@
$2\leq p,q\leq\infty,\quad\frac{2}{q}+\frac{2}{p}=1,\quad q\neq 2.$ (1.4)
@@ 2. $(ii)$
The analysis performed in [41] is based on the dimension $n\geq 2$ being odd or even, nevertheless using the strategy described in [41], an analogous result can also be obtained when $n=1$. In the following, we discuss this briefly for the hyperbolic NLS (1.1) with $k\geq 2$. First note that arguing as in [41, Proposition 2.1] one can obtain:
@@
$\|e^{it\Box}f\|_{L^{2k+2}_{t}W^{\frac{1}{2}-\frac{1}{k},\frac{2k+2}{k-1}}_{x}H^{\frac{1}{2}+\varepsilon}_{y}}+\Big\|\int_{0}^{t}e^{i(t-\tau)\Box}\prod^{2k+1}_{j=1}\widetilde{u}_{j}d\tau\Big\|_{L^{2k+2}_{t}W^{\frac{1}{2}-\frac{1}{k},\frac{2k+2}{k-1}}_{x}H^{\frac{1}{2}+\varepsilon}_{y}}$ (1.5)
@@
$\|e^{it\Box}P_{\leq N}\phi\|_{L^{p}_{t,\mathbf{x}}([0,1]\times\mathcal{M})}\lesssim N^{1-\frac{4}{p}}\|P_{\leq N}\phi\|_{L^{2}(\mathcal{M})}.$ (1.6)
@@
$\mathcal{T}:=\{t\in[0,1]:qN^{2}\big|t-a/q\big|\leq N^{2\sigma}\,\,\text{for some}\,\,q\leq N^{2\sigma},\,\text{and}\,(a,q)=1\},$ (1.7)
and kernel estimate due to Weyl’s method. In our situation, we introduce the short time intervals $I_{N}$ ((depending on dyadics $N)$ instead of the major arc $\mathcal{T}$; and complete the proof of Proposition 1.4 by an elementary argument that replaces the Weyl sums estimate. In particular, taking advantage of dispersive decay separately in both periodic ((thanks to [30, Theorem $5.3$])) and the real line directions, we provide a simple proof of Strichartz type estimate for the kernel over $I_{N}$, whereas the proof for analogous estimate [31, Proposition $2.3$] is much more involved. See Subsection 3.1 for the details.
@@
$\begin{split}\|e^{it\Box}P_{\leq N}\phi\|_{\ell^{q}L^{p}(\mathbb{R},L^{p}(\mathbb{R}\times\mathbb{T}))}&\lesssim N^{1-\frac{4}{p}}\|P_{\leq N}\phi\|_{L^{2}(\mathbb{R}\times\mathbb{T})}.\end{split}$ (1.8)
@@
$(i\partial_{t}+\partial_{x}\partial_{y})u=0,\quad x\in\mathbb{R},\,y\in\mathbb{T}.$ (1.9)
@@
$\|e^{it\partial_{x}\partial_{y}}P_{\leq N}f\|_{L^{4}_{t,x,y}([0,1]\times\mathbb{R}\times\mathbb{T})}\lesssim N^{\frac{1}{4}}\|f\|_{L^{2}(\mathbb{R}\times\mathbb{T})},$ (1.10)
which is observed in the context of Boltzmann equation, [5, Theorem 2.1]. Therefore, in view of the estimates (3.1) and (1.10), the linear part of (1.1) and the equation (1.9) behave very differently with respect to dispersive estimates. This is in contrast with the case $x\in\mathbb{R}$, $y\in\mathbb{R}$ when the linear HNLS in (1.1) and the equation (1.9) are equivalent. We also point out that when the spatial domain is $\mathbb{T}^{2}$, the $L^{4}$ Strichartz estimate for the linear group $e^{it\Box}$ exhibits the same loss of $N^{\frac{1}{4}}$ as in (1.10), see [43, 18].
@@ 1.1. Outline of the paper
In Section 2, we introduce notations and function spaces. In Section 3, we prove non-scale invariant local-in-time Strichartz estimates, and by an $\epsilon$-removal argument, we then obtain scale invariant versions of these Strichartz estimates. In Section 4, we upgrade the Strichartz estimates of Section 3 to the global Strichartz estimates. Lastly, in Section 5, we prove Theorem 1.1 by using the results in Section 4.
@@ 2.3. Function spaces
    $\|v\|_{V^{p}}:=\sup_{\{t_{k}\}_{k=0}^{K}\in\mathcal{Z}}\Big(\sum_{k=1}^{K}\|v(t_{k})-v(t_{k-1})\|_{\mathcal{H}}^{p}\Big)^{\frac{1}{p}}<\infty.$ (2.1)
@@ 3. Local Strichartz estimates
$\|e^{it\Box_{\mathbf{x}}}P_{\leq N}\phi\|_{L^{4}_{t,\mathbf{x}}([0,1]\times\mathcal{M})}\lesssim_{\epsilon}N^{\epsilon}\|P_{\leq N}\phi\|_{L^{2}(\mathcal{M})}.$ (3.1)
In [24], there is an alternative proof of the $L^{4}$ Strichartz estimate (3.1) based on decoupling estimates due to Guth-Maldague-Oh [19]. Note that interpolating the estimate (3.1) and the following straightforward estimate
@@ 3.1. The $\epsilon$-Removal Argument
$K_{N}(t,\mathbf{x})$ $=[e^{it\Box}P_{\leq N}\delta_{0}](\mathbf{x})$ (3.3)
@@
$\int_{0}^{N^{1-\delta}}$ $p\mu^{p-1}\big|\{(t,\mathbf{x})\in[0,1]\times\mathcal{M}:|(e^{it\Box}P_{\leq N}f)(\mathbf{x})|>\mu\}\big|d\mu$ (3.4)
@@
$\mu^{2}|\Omega_{\omega}|^{2}$ $\lesssim\langle e^{it\Box}P_{\leq N}f,\mathds{1}_{\Omega_{\omega}}\rangle_{L^{2}_{t,\mathbf{x}}}^{2}$ (3.5)
@@
$|K_{N}(t,\mathbf{x})|\lesssim\begin{cases}N|t|^{-\frac{1}{2}},\quad\text{for all}\,\,t\in\mathbb{R},\\ |t|^{-1},\qquad\text{for}\,\,|t|\leq N^{-1}.\end{cases}$ (3.6)
@@
$|K_{NI}(t,x)|\lesssim|t|^{-\frac{1}{2}},\quad\forall t\in\mathbb{R}.$ (3.7)
Thus, by (3.3), the estimate (3.7) and the straightforward bound $|K_{NS}|\lesssim N$ give rise to the first estimate in (3.6) which holds for all times. Then, due to the estimate $(5.9)$ in the proof of [30, Theorem $5.3$], we also get the bound
$|K_{NS}(t,y)|\lesssim|t|^{-\frac{1}{2}},\quad\text{for}\,\,|t|\leq N^{-1}.$ (3.8)
@@
$|\langle\mathds{1}_{\Omega_{\omega}},(K_{N}-K_{N,0})*\mathds{1}_{\Omega_{\omega}}\rangle_{L^{2}_{t,\mathbf{x}}}|$ $\lesssim\|\mathds{1}_{\Omega_{\omega}}\|_{L^{1}_{t,\mathbf{x}}}\|(K_{N}-K_{N,0})*\mathds{1}_{\Omega_{\omega}}\|_{L^{\infty}_{t,\mathbf{x}}}$ (3.9)
@@
Notice that, when $0<t<N^{-2}$, we have $\epsilon_{N}(t)=Nt^{-\frac{1}{2}}$, and for $N^{-2}\leq t\leq N^{-1}$, we have $\epsilon_{N}(t)=t^{-1}$. It is worth noting that though [31, Proposition $2.3$] is applicable in our analysis in the rest (when approximating kernel is defined over $\mathcal{T}$, see (1.7)), here, we present a straightforward proof thanks to the decay in Lemma 3.4.
@@
$|\langle\mathds{1}_{\Omega_{\omega}},K_{N,0}*\mathds{1}_{\Omega_{\omega}}\rangle_{L^{2}_{t,\mathbf{x}}}|$ $\lesssim\|\mathds{1}_{\Omega_{\omega}}\|_{L^{r^{\prime}}_{t,\mathbf{x}}}\|K_{N,0}*\mathds{1}_{\Omega_{\omega}}\|_{L^{r}_{t,\mathbf{x}}}$ (3.10)

### v3

@@ 1. Introduction
(1.1) $\begin{cases}i\partial_{t}u+\Box u=\pm|u|^{2k}u,\\ u(t,\mathbf{x})|_{t=0}=u_{0}(\mathbf{x}),\end{cases}\quad(t,\mathbf{x})\in\mathbb{R}\times\mathcal{M},$
@@
(1.3) $\|e^{it\partial_{x}\partial_{y}}f\|_{L^{q}_{t}L^{p}_{x,y}(\mathbb{R}\times\mathbb{R}^{2})}\lesssim\|f\|_{L^{2}(\mathbb{R}^{2})}$
@@
(1.4) $2\leq p,q\leq\infty,\quad\frac{2}{q}+\frac{2}{p}=1,\quad q\neq 2.$
@@ 2. $(ii)$
The analysis performed in [42] is based on the dimension $n\geq 2$ being odd or even, nevertheless using the strategy described in [42], an analogous result can also be obtained when $n=1$. In the following, we discuss this briefly for the hyperbolic NLS (1.1) with $k\geq 2$. First note that arguing as in [42, Proposition 2.1] one can obtain:
@@
(1.5) $\|e^{it\Box}f\|_{L^{2k+2}_{t}W^{\frac{1}{2}-\frac{1}{k},\frac{2k+2}{k-1}}_{x}H^{\frac{1}{2}+\varepsilon}_{y}}+\Big\|\int_{0}^{t}e^{i(t-\tau)\Box}\prod^{2k+1}_{j=1}\widetilde{u}_{j}d\tau\Big\|_{L^{2k+2}_{t}W^{\frac{1}{2}-\frac{1}{k},\frac{2k+2}{k-1}}_{x}H^{\frac{1}{2}+\varepsilon}_{y}}$
@@
(1.6) $\|e^{it\Box}P_{\leq N}\phi\|_{L^{p}_{t,\mathbf{x}}([0,1]\times\mathcal{M})}\lesssim N^{1-\frac{4}{p}}\|P_{\leq N}\phi\|_{L^{2}(\mathcal{M})}.$
@@
(1.7) $\mathcal{T}:=\{t\in[0,1]:qN^{2}\big|t-a/q\big|\leq N^{2\sigma}\,\,\text{for some}\,\,q\leq N^{2\sigma},\,\text{and}\,(a,q)=1\},$
and kernel estimate due to Weyl’s method. In our situation, we introduce the short time intervals $I_{N}$ ((depending on dyadics $N)$ instead of the major arc $\mathcal{T}$; and complete the proof of Proposition 1.4 by an elementary argument that replaces the Weyl sums estimate. In particular, taking advantage of dispersive decay separately in both periodic ((thanks to [31, Theorem $5.3$])) and the real line directions, we provide a simple proof of Strichartz type estimate for the kernel over $I_{N}$, whereas the proof for analogous estimate [32, Proposition $2.3$] is much more involved. See Subsection 3.1 for the details.
@@
(1.8) $\begin{split}\|e^{it\Box}P_{\leq N}\phi\|_{\ell^{q}L^{p}(\mathbb{R},L^{p}(\mathbb{R}\times\mathbb{T}))}&\lesssim N^{1-\frac{4}{p}}\|P_{\leq N}\phi\|_{L^{2}(\mathbb{R}\times\mathbb{T})}.\end{split}$
@@
(1.9) $(i\partial_{t}+\partial_{x}\partial_{y})u=0,\quad x\in\mathbb{R},\,y\in\mathbb{T}.$
@@
(1.10) $\|e^{it\partial_{x}\partial_{y}}P_{\leq N}f\|_{L^{4}_{t,x,y}([0,1]\times\mathbb{R}\times\mathbb{T})}\lesssim N^{\frac{1}{4}}\|f\|_{L^{2}(\mathbb{R}\times\mathbb{T})},$
which is observed in the context of the Boltzmann equation, [5, Theorem 2.1]. Therefore, in view of estimates (3.1) and (1.10), the linear part of (1.1) and the equation (1.9) behave very differently with respect to dispersive estimates. This is in contrast with the case $x\in\mathbb{R}$, $y\in\mathbb{R}$ when the linear HNLS in (1.1) and equation (1.9) are equivalent. We also point out that when the spatial domain is $\mathbb{T}^{2}$, the $L^{4}$ Strichartz estimate for the linear group $e^{it\Box}$ exhibits the same loss of $N^{\frac{1}{4}}$ as in (1.10), see [44, 19]. We have the following slight improvement of (1.10), which is of independent interest; for its proof, see Section 3.2.
Theorem 1.8.
Let $(x,y)\in\mathbb{R}\times\mathbb{T}$. Then, for $1\leq N$, we have
$\|e^{it\partial_{x}\partial_{y}}P_{\leq N}f\|_{L^{4}_{t,x,y}([0,1]\times\mathbb{R}\times\mathbb{T})}\lesssim(\log(N))^{\frac{1}{4}}\|f\|_{L^{2}(\mathbb{R}\times\mathbb{T})}+N^{\frac{1}{4}}\|\widehat{f}(\xi,n)\|_{l^{4}_{n}L^{2}_{\xi}}.$
The improvement comes from the fact that in the right hand-side of (1.10) we have $\|\widehat{f}(\xi,n)\|_{l^{2}_{n}L^{2}_{\xi}},$ i.e., we have an improvement in the $n$ summability condition.
@@ 1.1. Outline of the paper
In Section 2, we introduce notation and function spaces. In Section 3, we prove non-scale invariant local-in-time Strichartz estimates, and by an $\epsilon$-removal argument, we then obtain scale invariant versions of these Strichartz estimates. In Section 4, we upgrade the Strichartz estimates of Section 3 to the global Strichartz estimates. Lastly, in Section 5, we prove Theorem 1.1 by using the results in Section 4.
@@ 2.3. Function spaces
    (2.1) $\|v\|_{V^{p}}:=\sup_{\{t_{k}\}_{k=0}^{K}\in\mathcal{Z}}\Big(\sum_{k=1}^{K}\|v(t_{k})-v(t_{k-1})\|_{\mathcal{H}}^{p}\Big)^{\frac{1}{p}}<\infty.$
@@ 3. Local Strichartz estimates
(3.1) $\|e^{it\Box_{\mathbf{x}}}P_{\leq N}\phi\|_{L^{4}_{t,\mathbf{x}}([0,1]\times\mathcal{M})}\lesssim_{\epsilon}N^{\epsilon}\|P_{\leq N}\phi\|_{L^{2}(\mathcal{M})}.$
In [25], there is an alternative proof of the $L^{4}$ Strichartz estimate (3.1) based on decoupling estimates due to Guth-Maldague-Oh [20]. Moreover, it is worth noting that after our paper was published, Deng-Fan-Zhao [12] removed the $\epsilon$-loss in the estimate (3.1) by a proof based on the combination of a kernel decomposition method with measure estimates for semi-algebraic sets. Hence, using the sharp Strichartz estimate, they established the global well-posedness for the cubic HNLS (1.1) in $L^{2}$ with small data, which is left from Theorem 1.1. Note that interpolating the estimate (3.1) and the following straightforward estimate
@@ 3.1. The $\epsilon$-Removal Argument
(3.3) $K_{N}(t,\mathbf{x})$ $=[e^{it\Box}P_{\leq N}\delta_{0}](\mathbf{x})$
@@
(3.4) $\int_{0}^{N^{1-\delta}}$ $p\mu^{p-1}\big|\{(t,\mathbf{x})\in[0,1]\times\mathcal{M}:|(e^{it\Box}P_{\leq N}f)(\mathbf{x})|>\mu\}\big|d\mu$
@@
(3.5) $\mu^{2}|\Omega_{\omega}|^{2}$ $\lesssim\langle e^{it\Box}P_{\leq N}f,\mathds{1}_{\Omega_{\omega}}\rangle_{L^{2}_{t,\mathbf{x}}}^{2}$
@@
(3.6) $|K_{N}(t,\mathbf{x})|\lesssim\begin{cases}N|t|^{-\frac{1}{2}},\quad\text{for all}\,\,t\in\mathbb{R},\\ |t|^{-1},\qquad\text{for}\,\,|t|\leq N^{-1}.\end{cases}$
@@
(3.7) $|K_{NI}(t,x)|\lesssim|t|^{-\frac{1}{2}},\quad\forall t\in\mathbb{R}.$
Thus, by (3.3), the estimate (3.7) and the straightforward bound $|K_{NS}|\lesssim N$ give rise to the first estimate in (3.6) which holds for all times. Then, due to the estimate $(5.9)$ in the proof of [31, Theorem $5.3$], we also get the bound
(3.8) $|K_{NS}(t,y)|\lesssim|t|^{-\frac{1}{2}},\quad\text{for}\,\,|t|\leq N^{-1}.$
@@
(3.9) $|\langle\mathds{1}_{\Omega_{\omega}},(K_{N}-K_{N,0})*\mathds{1}_{\Omega_{\omega}}\rangle_{L^{2}_{t,\mathbf{x}}}|$ $\lesssim\|\mathds{1}_{\Omega_{\omega}}\|_{L^{1}_{t,\mathbf{x}}}\|(K_{N}-K_{N,0})*\mathds{1}_{\Omega_{\omega}}\|_{L^{\infty}_{t,\mathbf{x}}}$
@@
Notice that, when $0<t<N^{-2}$, we have $\epsilon_{N}(t)=Nt^{-\frac{1}{2}}$, and for $N^{-2}\leq t\leq N^{-1}$, we have $\epsilon_{N}(t)=t^{-1}$. It is worth noting that though [32, Proposition $2.3$] is applicable in our analysis in the rest (when approximating kernel is defined over $\mathcal{T}$, see (1.7)), here, we present a straightforward proof thanks to the decay in Lemma 3.4.
@@
(3.10) $|\langle\mathds{1}_{\Omega_{\omega}},K_{N,0}*\mathds{1}_{\Omega_{\omega}}\rangle_{L^{2}_{t,\mathbf{x}}}|$ $\lesssim\|\mathds{1}_{\Omega_{\omega}}\|_{L^{r^{\prime}}_{t,\mathbf{x}}}\|K_{N,0}*\mathds{1}_{\Omega_{\omega}}\|_{L^{r}_{t,\mathbf{x}}}$
@@
3.2. Proof of Theorem 1.8
Set $u(t,x,y)=e^{it\partial_{x}\partial_{y}}P_{\leq N}f\,.$ We have
$u^{2}(t,x,y)=\sum_{n_{1},n_{2}\in\mathbb{Z}}\,\,\int_{\mathbb{R}^{2}}e^{2\pi i((\xi_{1}n_{1}+\xi_{2}n_{2})t+(\xi_{1}+\xi_{2})x+(n_{1}+n_{2})y)}\\ \varphi(\xi_{1}/N)\varphi(n_{1}/N)\,\varphi(\xi_{2}/N)\varphi(n_{2}/N)\,\widehat{f}(\xi_{1},n_{1})\widehat{f}(\xi_{2},n_{2})d\xi_{1}d\xi_{2}.$
Write
$u^{2}(t,x,y)=u_{1}(t,x,y)+u_{2}(t,x,y),$
where
$u_{1}(t,x,y)=\sum_{n_{1}\neq n_{2}}\,\,\int_{\mathbb{R}^{2}}e^{2\pi i((\xi_{1}n_{1}+\xi_{2}n_{2})t+(\xi_{1}+\xi_{2})x+(n_{1}+n_{2})y)}\\ \varphi(\xi_{1}/N)\varphi(n_{1}/N)\,\varphi(\xi_{2}/N)\varphi(n_{2}/N)\,\widehat{f}(\xi_{1},n_{1})\widehat{f}(\xi_{2},n_{2})d\xi_{1}d\xi_{2}$
and
$u_{2}(t,x,y)=\sum_{n\in\mathbb{Z}}\,\,\int_{\mathbb{R}^{2}}e^{2\pi i((\xi_{1}+\xi_{2})(x+nt)+2ny)}\varphi(\xi_{1}/N)\,\varphi(\xi_{2}/N)\varphi^{2}(n/N)\,\widehat{f}(\xi_{1},n)\widehat{f}(\xi_{2},n)d\xi_{1}d\xi_{2}.$
We have that
$\|u_{2}(t,x,\cdot)\|^{2}_{L^{2}(\mathbb{T})}=\\ \sum_{n\in\mathbb{Z}}\,\Big|\int_{\mathbb{R}^{2}}e^{2\pi i((\xi_{1}+\xi_{2})(x+nt))}\varphi(\xi_{1}/N)\,\varphi(\xi_{2}/N)\varphi^{2}(n/N)\,\widehat{f}(\xi_{1},n)\widehat{f}(\xi_{2},n)d\xi_{1}d\xi_{2}\Big|^{2}\,.$
Set
$F_{n}(t,x)=\int_{\mathbb{R}^{2}}e^{2\pi i((\xi_{1}+\xi_{2})(x+nt))}\varphi(\xi_{1}/N)\,\varphi(\xi_{2}/N)\varphi^{2}(n/N)\,\widehat{f}(\xi_{1},n)\widehat{f}(\xi_{2},n)d\xi_{1}d\xi_{2}.$
We need to evaluate $\|F_{n}\|_{L^{2}([0,1]\times\mathbb{R})}.$ Write
$F_{n}(t,x)=\int_{\mathbb{R}}e^{2\pi i\xi(x+nt)}\Big(\int_{\mathbb{R}}\varphi(\xi_{1}/N)\,\varphi((\xi-\xi_{1})/N)\varphi^{2}(n/N)\,\widehat{f}(\xi_{1},n)\widehat{f}(\xi-\xi_{1},n)d\xi_{1}\Big)d\xi.$
Using the variable change $x^{\prime}=x+nt$ (the $t$ integration disappears), we can write
$\|F_{n}\|_{L^{2}([0,1]\times\mathbb{R})}^{2}=\int_{\mathbb{R}}\Big|\int_{\mathbb{R}}\varphi(\xi_{1}/N)\,\varphi((\xi-\xi_{1})/N)\varphi^{2}(n/N)\,\widehat{f}(\xi_{1},n)\widehat{f}(\xi-\xi_{1},n)d\xi_{1}\Big|^{2}d\xi.$
Using Cauchy-Schwarz in the $\xi_{1}$ integration we get
$\|F_{n}\|_{L^{2}([0,1]\times\mathbb{R})}^{2}\lesssim N\|\widehat{f}(\xi,n)\|_{L^{2}_{\xi}}^{4}$
which in turn implies
$\|u_{2}\|^{2}_{L^{2}_{t,x,y}([0,1]\times\mathbb{R}\times\mathbb{T})}\lesssim N\|\widehat{f}(\xi,n)\|^{4}_{l^{4}_{n}L^{2}_{\xi}}.$
Let us next estimate $\|u_{1}\|_{L^{2}_{t,x,y}([0,1]\times\mathbb{R}\times\mathbb{T})}.$ After some classical reductions based on the Littlewood-Paley theory and the use of the Cauchy-Schwarz inequality, it suffices to bound
${\rm mes}(A_{N}(\tau,\xi,n)),$
where
$A_{N}(\tau,\xi,n)=\big\{(\xi_{1},n_{1})\in\mathbb{R}\times\mathbb{Z}\,:\,|\tau+\xi_{1}n_{1}+(\xi-\xi_{1})(n-n_{1})|\leq C,\,\,n_{1}\neq n-n_{1}\\ |n_{1}|\leq 2N,\,\,\,|n-n_{1}|\leq 2N,\,\,\,|\xi_{1}|\leq 2N,\,\,\,|\xi-\xi_{1}|\leq 2N\big\},$
and where $\tau,\xi\in\mathbb{R}$, $n\in\mathbb{Z}$ and $N$ is a dyadic integer. We have
$\xi_{1}n_{1}+(\xi-\xi_{1})(n-n_{1})=(2n_{1}-n)\xi_{1}+\xi n-\xi n_{1}.$
The key point is that for $(\xi_{1},n_{1})\in A_{N}(\tau,\xi,n)$ we have that $2n_{1}-n\neq 0$. Also, we have that if $A_{N}(\tau,\xi,n)$ is nonempty then $|n|\leq 4N$. For each fixed $n_{1}\in A_{N}(\tau,\xi,n)$ the measure of the possible $\xi_{1}$ is bounded by
$\frac{C}{2n_{1}-n}$
and therefore
${\rm mes}(A_{N}(\tau,\xi,n))\lesssim\sum_{|n_{1}|\leq 2N,2n_{1}\neq n}\frac{1}{2n_{1}-n},$
where $n$ is an integer in the interval [4N,4N][-4N,4N]. Therefore
${\rm mes}(A_{N}(\tau,\xi,n))\lesssim\sum_{n_{1}=1}^{10N}\frac{1}{n_{1}}\lesssim\log(N)$
which implies that
$\|u_{1}\|_{L^{2}_{t,x,y}([0,1]\times\mathbb{R}\times\mathbb{T})}^{2}\lesssim\log(N)\|f\|_{L^{2}(\mathbb{R}\times\mathbb{T})}^{4}\,.$
This completes the proof of Theorem 1.8.